Design a rigid coupling is used to connect a 45 kW,1440 rpm electric motor to a 

centrifugalpump. The starting torque of the motor is225% of the rated torque


1
Expert's answer
2021-04-21T07:05:42-0400

power "P = 45KW = 45,000 Nm"


RPM N = 1440


"T = \frac{P \times 60}{2\pi N} = \frac{45000 \times 60}{2\pi 1440} = 298.41 Nm"


As starting torque of the motor is 225% of rated torque.


"T_{max} = 2.25 \times T = 2.25 \times 298.41 = 671.42 Nm"